In the mid-1990s, one of the main debates in Major League Baseball was whether Derek Jeter or David Justice was a better baseball player. One key metric in this debate was their respective batting averages, calculated as the number of hits divided by number of at-bats. The following information is available regarding their batting averages in the years 1995 and 1996:

1995

1996

Combined

Player

Number of Hits

Number of At-Bats

Number of Hits

Number of At-Bats

Number of Hits

Number of At-Bats

Derek Jeter

12

48

183

582

195

630

David Justice

104

411

45

140

149

551

  1. Looking at the years 1995 and 1996 individually, who has a higher batting average?
  1. Look at the two years combined, who has a higher batting average?
  1. Compare your answers to parts a and b. What phenomena is at play here, and why does it occur?
Answer #1

Requireenta: Year Derek jeter 1995 12/48 = 0.250 1996 183/582 = 0.314 David justice 104/411 = 0.253 45/140 = 0.321 The above

Requirement c: If we compare the answers to parts a and b, there was a contrasting result. Eventhough david justice had highe